# MUSE-Net: Missingness-aware mUlti-branching Self-attention Encoder for Irregular Longitudinal Electronic Health Records

### v1

@@ Abstract
Index Terms: Irregularly spaced time series, Multivariate longitudinal records, Data imputation, Imbalanced dataset, Multi-task Gaussian process, Self-attention encoder
1 Introduction
@@
However, data mining of longitudinal EHRs poses distinctive research challenges due to the observational nature of EHRs. Unlike well-defined, randomized longitudinal experiments in clinical trials that collect data on a fixed schedule and ensure high data quality, EHRs are recorded only when patients receive care or doctors provide services. The information collected and the timing of its collection are not determined by researchers, resulting in EHR data that are often complex and highly heterogeneous [11]. Specifically, fully utilizing real-world EHRs for reliable data-driven decision-making calls upon addressing several challenges:
@@
(2) Incomplete datasets and imbalanced class distributions. Real-world EHRs suffer from the issues of missing values and imbalanced data. Due to the nature of clinical practice, not all information is recorded for every patient visit, leading to incomplete datasets. Additionally, EHR data often exhibit a significantly imbalanced distribution, with certain health outcomes or characteristics being underrepresented compared to more common ones. For instance, rare diseases or adverse drug reactions [12, 13] may have very few instances compared to more prevalent conditions. In the existing literature, a wide array of statistical and machine learning techniques have been designed to tackle the missing value and imbalanced data issues [14, 15], which, however, are less applicable in the context of modeling irregular longitudinal EHRs. The presence of missing values and imbalanced class distributions will introduce further difficulties in effective model training using longitudinal EHRs, leading to biased or inaccurate predictions if not properly addressed.
@@ 2 Research Background
The integration of data-driven modeling methodologies and EHR data has transformed the healthcare field, which provides unprecedented opportunities for clinical decision support [8, 16, 17]. Extensive research has been conducted to develop data-driven models for disease identification using non-longitudinal EHRs that consist of static or cross-sectional patient health information [18, 19]. For example, Huang et al. [20] employed naive Bayes, decision tree, and nearest neighbor algorithms incorporating feature selection methods to determine key factors affecting type II diabetes control and further identify individuals who exhibit suboptimal diabetes control status. Hong et al. [21] developed a multi-class classification method to analyze clinical data in identifying patients with obesity and various comorbidities using logistic regression, support vector machine, and decision tree algorithms. A comprehensive review on machine learning of non-longitudinal EHRs can be referred to [11, 22]. However, those traditional methods are limited in capturing temporal patterns in patient health status over time. This limitation can result in less accurate predictions for conditions that are heavily dependent on longitudinal health trajectories [18, 23]. Moreover, traditional machine learning approaches often require manual feature engineering, which is not only time-consuming and prone to error but also demands significant domain expertise.
@@
With the growing availability of longitudinal EHRs, increasing interests have been devoted to developing advanced data-driven models to capture critical temporal information for disease prediction. Owing to the outstanding performance and strong capability in pattern recognition, deep learning has been widely explored to mine complexly structured data leveraging the nonlinear interactions among neurons within and across various hidden layers [24, 25, 26, 5]. Advanced network architectures have been crafted for modeling time series or sequential data. For example, recurrent neural networks (RNNs) including long short-term memory (LSTM) are among the most commonly used deep learning models to analyze medical time series for various clinical tasks [27, 28, 29]. Additionally, temporal convolutional networks (TCNs) have been recognized as a robust alternative to RNNs for modeling longitudinal clinical signals [7, 6]. The unique architecture of dilated and causal 1-D convolution and residual blocks [30] allows TCNs to access a broader range of sequential information for capturing long-term temporal dependencies, thereby outperforming RNNs in various tasks involving mining temporal data [7, 6, 31]. However, most existing RNNs and TCNs are designed with the assumption that the longitudinal records are collected at a constant rate and require the neighboring samples to appear at fixed distances to facilitate the convolution or recurrent operations. This assumption is not valid in many real-world databases, making traditional RNNs or TCNs less effective in modeling irregular-spaced longitudinal EHRs.
@@
To cope with the issue of irregular time intervals, many modified RNN architectures have been developed. For example, the time-aware LSTM (TLSTM) was designed to account for non-uniform sampling intervals through a time decay mechanism [32]. Che et al. [28] also implemented a decay mechanism and proposed a GRU-D model for time series signals, allowing the network to better capture temporal dependencies even when data points were incomplete or irregularly sampled. A comprehensive review of modified RNN architectures for irregular sampled time series can be found in [33, 34]. However, most existing time-aware RNNs assume that the impact of historical risk factors on disease prediction proportionally decays over time, which may not be true in describing complex disease trajectories. Additionally, RNNs have been widely recognized as computationally inefficient for modeling large-scale, long sequential data due to their sequential processing nature.
@@
The self-attention encoder, a cornerstone of the transformer architecture [35], has revolutionized the field of natural language processing (NLP) with its unique advantages. Unlike RNNs and TCNs that process data sequentially and require uniform sampling intervals, the self-attention mechanism allows the model to weigh the importance of different parts of the input sequences relative to each other. This feature enables parallel processing and enhances the model’s ability to capture long-range dependencies in irregularly spaced longitudinal signals. For instance, Li et al. proposed to transform time series data into images and adopted the Vision Transformer to model irregularly sampled time series signals [36]. Tipirneni and Reddy developed a method that integrated a continuous value embedding technique with the self-attention architecture to model irregularly sampled clinical time series by treating them as a set of observation triplets (consisting of time, variable, and value)[37]. Huang et al. developed a Deformable Neighborhood Attention Transformer (DNA-T) to capture local and global dependencies in medical time series data through deformable attention mechanisms[38]. Those self-attention-based models demonstrate significant improvements over traditional methods, particularly in handling the irregularities inherent in real-world longitudinal datasets. Despite the strengths of self-attention encoders/transformer architecture, it is worth noting that most standard deep learning algorithms are designed under the assumption that data is ideal, which cannot be directly applied to mine longitudinal EHR databases with missing values and imbalanced class distributions. Little has been done to design an effective framework to model real-world longitudinal EHRs by accounting for the issues of missing values, imbalanced data distribution, and irregular time intervals for robust decision-making.
3 Research Methodology
@@
Figure 1: The proposed MUSE-Net framework: (a) Multi-task Gaussian process for data imputation, and missing value masks generation; The imputed data and missing value masks are processed by the proposed MUSE-Net, which consist of (b) Missingness-aware Self-attention Encoder and (c) Multi-branching output layer.
3.1 Multi-task Gaussian Process (MGP) for Missing Value Imputation
A Gaussian process (GP) is a flexible non-parametric Bayesian model where any collection of random variables follows a joint Gaussian distribution [39, 40, 41], which have been widely used to model complex time series data [42, 43]. GP-based temporal models provide a way to determine the distribution of a variable at any arbitrary point in time, making them intrinsically capable of dealing with missing value imputations that involve irregularly spaced longitudinal data. Note that single-task GPs are limited in their ability to model correlations across multiple related tasks (i.e., different medical variables), making them less effective for modeling multi-variate EHR data. To account for the multi-variate nature, we adopt a multi-task GP (MGP) [44], which is an extension to single-task GPs, to capture both variable interactions and temporal correlations for imputing missing values in irregular multi-variate longitudinal EHRs.
@@ 3.1 Multi-task Gaussian Process (MGP) for Missing Value Imputation
where $y_{pm}(t)$ is the observed value of variable $m$ for patient $p$ at time $t$, $\sigma^{2}_{m}$ is the noise term for the $m$th task (variable), $\mathcal{K}_{M}(m,m^{{}^{\prime}})$ captures the similarities between tasks, and $\mathcal{K}_{t}(t,t^{{}^{\prime}})$ is a temporal correlation function. Note that a fundamental property of time series data is that observations at neighboring points tend to be similar. As such, we use the squared exponential kernel function $\mathcal{K}_{t}(t,t^{{}^{\prime}})=\text{exp}(-\frac{(t-t^{{}^{\prime}})^{2}}{2\theta^{2}})$ with a lengthscale parameter $\theta$. Hence, the prior distribution for the fully observed multivariate longitudinal records, $\boldsymbol{y}^{o}_{p}$, can be represented by:
@@
Finally, the missing values in $Y_{p}$ are estimated as $\hat{\boldsymbol{y}}_{p}^{u}=\mu_{\boldsymbol{y}_{p}^{u}}$, which are combined with the observed values $\boldsymbol{y}_{p}^{o}$ to form an imputed dataset $\hat{\mathcal{D}}=\{(\boldsymbol{t}_{p},\hat{Y}_{p},l_{p})\}_{p=1}^{N}=\{((\boldsymbol{t}_{p}^{o},\text{Vect}^{-1}(\boldsymbol{y}_{p}^{o})),(\boldsymbol{t}_{p}^{u},\text{Vect}^{-1}(\hat{\boldsymbol{y}}_{p}^{u})),l_{p})\}_{p=1}^{N}$, where $\text{Vect}^{-1}(\cdot)$ is the inverse operation of $\text{Vect}(\cdot)$.
@@
In addition, to accelerate MGP Imputation, we adopt the black box matrix-matrix multiplication framework [45] to integrate the modified preconditioned conjugate gradient (mPCG) approach with GPU acceleration into the data imputation workflow. Please refer to Appendix 8.3 for the detailed procedure for GPU acceleration of MGP imputation.
3.2 Missingness-aware Multi-branching Self-Attention Encoder (MUSE-Net)
@@
We further propose a novel missingness-aware multi-branching self-attention encoder network (MUSE-Net) to process the imputed longitudinal EHRs. This model is designed to not only effectively capture crucial temporal correlations in multi-variate irregular longitudinal data but also recognize missingness patterns and account for the imbalanced data issue for enhanced classification performance. As shown in Fig. 1(b) and (c), this model consists of 3 key modules: time-aware self-attention encoder, missing value masks, and multi-branching outputs.
3.2.1 Time-aware Self-attention Encoder
@@
We propose to adapt the traditional self-attention encoder [35] to incorporate the information of elapsed time between consecutive records in irregular longitudinal EHRs. Specifically, there are three building blocks in our time-aware self-attention encoder:
@@
Elapsed time-based positional encoding: The time order of the input longitudinal sequence plays a crucial role in time series analysis. However, this information is often ignored in traditional attention-based encoders because it does not incorporate any recurrent or convolution operations. To address this issue, an elapsed time-based positional encoding is added into the input sequence at the beginning of the self-attention encoder as shown in Fig. 1(b), such that the temporal information regarding the irregularly sampled EHRs is incorporated into the model. In our study, we employ the following sinusoidal encoding:
$PE_{t_{p,i},2m}=\text{sin}(t_{p,i}/10000^{2m/M})\quad\quad\quad PE_{t_{p,i},2m+1}=\text{cos}(t_{p,i}/10000^{2m/M})$ (5)
@@
where $t_{p,i}$ is the time position of each observation to account for the irregularly spaced time interval, and $2m$ and $12m+$ 1 represent the ($2m$)-th and ($12m+$ 1)-th variable dimensions, respectively. The sinusoid prevents positional encodings from becoming too large, introducing extra difficulties in network optimization. The embedded elapsed time-based positional features will then be combined with the imputed longitudinal sequence to generate the input of the self-attention module: $X_{p}=PE_{p}+\hat{Y}_{p}$, where $PE_{p}\in\mathbb{R}^{T_{p}\times M}$ is a matrix with elements defined in Eq. (5).
@@
Self-attention module: The self-attention module is the fundamental mechanism that allows the network to learn dependencies in the longitudinal data. It maps a query set and a set of key-value pairs to an output. In a typical single-head self-attention, the key, query, and value matrices, denoted as $K\in\mathbb{R}^{T_{p}\times M}$, $Q\in\mathbb{R}^{T_{p}\times M}$, and $V\in\mathbb{R}^{T_{p}\times M}$, are computed by taking the input $X_{p}$:
@@
where $\text{softmax}(\boldsymbol{x})_{i}=\frac{e^{x_{i}}}{\sum_{j}e^{x_{j}}}$ outputs the weight assigned to each element in $V$. Note that the dot product in the softmax function is scaled down by $\sqrt{M}$. This is essential to prevent the dot product values from growing too large, especially when the dimensionality of $\hat{Y}_{p}$ or $X_{p}$ is large, introducing instabilities in the training process. The self-attention mechanism enables the network to access all the information of the input with the flexibility of focusing on certain important elements over the entire sequence, unlike the RNNs or CNNs which focus more on the neighboring elements.
The multi-head attention module is an extension of single-head self-attention to capture different relations among multivariate time series, which is adopted in our study to model complex longitudinal EHRs. Specifically, we define multiple matrices for the keys, queries, and values by applying multiple network operations, $\mathcal{F}_{K}^{l}(\cdot)$’s, $\mathcal{F}_{Q}^{l}(\cdot)$’s, and $\mathcal{F}_{V}^{l}(\cdot)$’s, to the input:
$K_{l}=\mathcal{F}_{K}^{l}(X_{p})\in\mathbb{R}^{T_{p}\times\frac{M}{h}},~{}~{}Q_{l}=\mathcal{F}_{Q}^{l}(X_{p})\in\mathbb{R}^{T_{p}\times\frac{M}{h}},~{}~{}V_{l}=\mathcal{F}_{V}^{l}(X_{p})\in\mathbb{R}^{T_{p}\times\frac{M}{h}}$ (8)
where $h$ is the number of attention heads and $l\in\{1,\dots,h\}$. The output of each attention head, $\text{Output}_{l}\in\mathbb{R}^{T_{p}\times\frac{M}{h}}$, is calculated by:
$\text{Output}_{l}=\text{softmax}(\frac{Q_{l}K_{l}^{T}}{\sqrt{M/h}})\times V_{l}$ (9)
@@
This multi-head attention module enables the network to jointly attend to information from different subspaces of multivariate time series at different time points. Simple visualization of the computing procedure of multi-head attention can be found in Fig. 2.
Figure 2: Illustration of multi-head attention with 2 heads. (a) Derivation of $Q,K,V$ matrices for each head with linear transformations of $X_{p}$: $Q_{l}=\mathcal{F}_{Q}^{l}(X_{p})=X_{p}W_{l}^{Q}$, $K_{l}=\mathcal{F}_{K}^{l}(X_{p})=X_{p}W_{l}^{K}$, $V_{l}=\mathcal{F}_{V}^{l}(X_{p})=X_{p}W_{l}^{V}$, $l=\{1,2\}$; (b) Compute key-query pairs for each head; (c) Compute attention weights from softmax function and further compute the output of heads 1 and 2 as the weighted sum of values; (d) The final output is the multiplication of the concatenation of all head outputs and another weight matrix $W^{O}$.
@@
3.2.2 Missing Value Masks
To further account for potential discrepancies between imputed values and actual observations, the missing value masks (see Fig. 1(a)) are incorporated as an additional input to the proposed model. We define the missing value masks for patient $p$ as $\tilde{X}_{p}$ with $[\tilde{X}_{p}]_{t,m}=1$ if $[Y_{p}]_{t,m}$ is missing; otherwise $[\tilde{X}_{p}]_{t,m}=0$. The masks will be processed by an independent time-aware self-attention encoder in parallel with the imputed sequences as shown in Fig. 1(b). This missingness-aware design enables the network to capture the relationship between non-missing and missing values. As a result, the network is capable of mitigating the effect of potential imputation errors during the MGP imputation process, thereby enhancing the overall predictive performance of the model.
3.2.3 Multi-branching Outputs
The self-attention encoder serves as a powerful feature extractor from irregular longitudinal EHRs for downstream tasks, such as binary classification. The classifier network, typically with a fully connected layer, is added after the feature extractor to interpret the abstracted features and make a final classification of the original input. The imbalanced data issue, a common problem in healthcare databases, incurs a more pronounced and direct negative impact on the classifier than on the feature extractor network. This is due to the fact that the classifier network directly interacts with the labels and is heavily influenced by the majority class during the training process [49]. This leads to a bias towards the majority class, visibly affecting the model’s performance. As such, careful design of the classifier network is needed to cope with the imbalanced data issue.

### v2

@@ Abstract
Note to Practitioners
This article is motivated by the growing need for robust machine learning models capable of handling the complexities of real-world EHRs, including irregular time intervals, missing data, and class imbalance. The proposed MUSE-Net model integrates advanced imputation via multi-task Gaussian processes with missingness masks, a time-aware self-attention encoder, and a multi-branching framework to enhance predictive accuracy and robustness. Additionally, MUSE-Net leverages an interpretable multi-head attention mechanism to provide transparent decision-making, allowing clinicians to trace model predictions back to key time points. This framework offers a practical and trustworthy solution for data-driven disease prediction and clinical decision support.
Index Terms:
Irregularly spaced time series, Multivariate longitudinal records, Data imputation, Imbalanced dataset, Multi-task Gaussian process, Self-attention encoder, Interpretable multi-head attention
I Introduction
@@
However, data mining of longitudinal EHRs poses distinct challenges due to the observational nature of EHRs. Unlike well-defined, randomized experiments in clinical trials that collect data on a fixed schedule and ensure high data quality, EHRs are recorded only when patients receive care or doctors provide services. The information collected and the timing of its collection are not determined by researchers, resulting in EHRs that are highly heterogeneous [11] and further introducing the following challenges in data-driven decision-making:
@@
(2) Incomplete data and imbalanced class distributions. EHRs suffer from the issues of missing values and imbalanced data. Due to the nature of clinical practice, not all information is recorded for every patient visit, leading to incomplete datasets. Additionally, EHR data often exhibit a significantly imbalanced distribution, with certain health outcomes or characteristics being underrepresented. For instance, rare diseases or adverse drug reactions [12, 13] may have very few instances compared to more prevalent conditions. In the existing literature, a wide array of statistical and machine learning techniques have been designed to tackle the missing value and imbalanced data issues [14, 15], which, however, are less applicable in the context of modeling irregular longitudinal EHRs. The presence of missing values and imbalanced class distributions will introduce further difficulties in effective model training using longitudinal EHRs, leading to biased or inaccurate predictions if not properly addressed.
@@ II Research Background
The integration of data-driven modeling and EHRs has transformed the healthcare field, which provides unprecedented opportunities for clinical decision support [8, 17, 18]. Extensive research has been conducted to develop data-driven models using non-longitudinal EHRs that consist of static or cross-sectional health information [19, 20]. For example, Huang et al. [21] employed naive Bayes, decision tree, and nearest neighbor algorithms incorporating feature selection methods to determine key factors affecting type II diabetes control and identify individuals who exhibit suboptimal diabetes control status. Hong et al. [22] developed a multi-class classification method to analyze clinical data in identifying patients with obesity and various comorbidities using logistic regression, support vector machine, and decision tree. A comprehensive review on machine learning of non-longitudinal EHRs can be referred to [11, 23]. However, those methods are limited in capturing temporal patterns in health status over time. This limitation can result in less accurate predictions for conditions that are heavily dependent on longitudinal health trajectories [19, 24]. Moreover, traditional machine learning approaches often require manual feature engineering, which is time-consuming and prone to error.
@@
With the growing availability of longitudinal EHRs, increasing interests have been devoted to developing advanced models to capture temporal information for disease prediction. Owing to the strong capability in pattern recognition, deep learning has been widely explored to mine complexly structured data [25, 26, 27, 5]. Advanced network architectures have been crafted for modeling time series or sequential data. For example, RNNs including long short-term memory (LSTM) are among the most commonly used models to analyze medical time series for various clinical tasks [28, 29, 30]. Additionally, temporal convolutional networks (TCNs) have been recognized as a robust alternative to RNNs for modeling longitudinal signals [7, 6, 31]. However, traditional RNNs and TCNs are designed with the assumption that the records are collected at a constant rate and require the neighboring samples to appear at fixed distances to facilitate the convolution or recurrent operations. This assumption is not valid in many real-world databases, making traditional RNNs or TCNs less effective in modeling irregular-spaced longitudinal EHRs.
@@
To cope with the issue of irregular time intervals, many modified RNN architectures have been developed. For example, the time-aware LSTM (TLSTM) was designed to account for non-uniform sampling intervals through a time decay mechanism [32]. Che et al. [29] also implemented a decay mechanism and proposed a GRU-D model, allowing the network to better capture temporal dependencies even when data points were incomplete or irregularly sampled. A comprehensive review of modified RNN architectures for irregular sampled time series can be found in [33, 34]. However, most existing time-aware RNNs assume that the impact of historical risk factors on disease prediction proportionally decays over time, which may not be true in describing complex disease trajectories. Additionally, RNNs have been widely recognized as computationally inefficient for modeling large-scale, long sequential data due to their sequential processing nature.
@@
The self-attention encoder, a cornerstone of the transformer architecture [35], has revolutionized the field of natural language processing (NLP). Unlike RNNs and TCNs that process data sequentially and require uniform sampling intervals, the self-attention mechanism allows the model to weigh the importance of different parts of the input sequences relative to each other. This feature enables parallel processing and enhances the model’s ability to capture long-range dependencies in irregular longitudinal signals. For instance, Li et al. proposed to transform time series data into images and adopted the Vision Transformer to model irregularly sampled time series signals [36]. Tipirneni and Reddy developed a method that integrated a continuous value embedding technique with self-attention to model irregularly sampled clinical time series by treating them as a set of observation triplets (time, variable, and value) [37]. Huang et al. developed a Deformable Neighborhood Attention Transformer to capture local and global dependencies in medical time series data through deformable attention mechanisms [38]. Manzini et al. introduced the Diabetic Attention with Relative representation Encoder for Type 2 Diabetes patients [39]. Those self-attention models demonstrate significant improvements over traditional methods, particularly in handling irregular longitudinal datasets.
Despite the strengths of self-attention encoders, most deep learning models assume well-structured data, making them less effective for longitudinal EHRs that are often incomplete, imbalanced, and irregularly sampled. Existing methods rarely integrate strategies to handle missing values, imbalanced class distributions, and temporal irregularity simultaneously, limiting their reliability in clinical applications. There remains a critical need for a framework that not only addresses these challenges holistically but also ensures robust and interpretable decision-making for real-world healthcare prediction tasks.
Figure 1: (a) MGP for data imputation and missing value masks generation; The imputed data and missing value masks are processed by MUSE-Net, which consist of (b) Missingness-aware Self-attention Encoder with Interpretable Multi-head Attention and (c) Multi-branching output layer.
III Research Methodology
@@
III-A MGP for Missing Value Imputation
A Gaussian process (GP) is a flexible non-parametric Bayesian model where any collection of random variables follows a joint Gaussian distribution [40, 41, 42], which have been widely used to model complex time series data [43, 44]. GP-based temporal models provide a way to determine the distribution of a variable at any arbitrary point in time, making them intrinsically capable of dealing with missing value imputations that involve irregularly spaced longitudinal data. Note that single-task GPs are limited in their ability to model correlations across multiple related tasks (i.e., different medical variables), making them less effective for modeling multi-variate EHR data. To account for the multi-variate nature, we adopt a multi-task GP (MGP) [45] to capture both variable interactions and temporal correlations for imputing missing values in irregular multi-variate longitudinal EHRs.
@@ III-A MGP for Missing Value Imputation
where $y_{pm}(t)$ is the observed value of variable $m$ for patient $p$ at time $t$, $\sigma^{2}_{m}$ is the noise term for the $m$th task (variable), $\mathcal{K}_{M}(m,m^{{}^{\prime}})$ captures the similarities between tasks, and $\mathcal{K}_{t}(t,t^{{}^{\prime}})$ is a temporal correlation function, which is defined as $\mathcal{K}_{t}(t,t^{{}^{\prime}})=\text{exp}(-\frac{(t-t^{{}^{\prime}})^{2}}{2\theta^{2}})$ with a lengthscale parameter $\theta$. Hence, the prior distribution for the fully observed multivariate longitudinal records, $\boldsymbol{y}^{o}_{p}$, can be represented by:
@@
Finally, the missing values in $Y_{p}$ are estimated as $\hat{\boldsymbol{y}}_{p}^{u}=\mu_{\boldsymbol{y}_{p}^{u}}$, which are combined with the observed values $\boldsymbol{y}_{p}^{o}$ to form an imputed dataset $\hat{\mathcal{D}}=\{(\boldsymbol{t}_{p},\hat{Y}_{p},l_{p})\}_{p=1}^{N}=\{((\boldsymbol{t}_{p}^{o},\text{Vect}^{-1}(\boldsymbol{y}_{p}^{o})),(\boldsymbol{t}_{p}^{u},\text{Vect}^{-1}(\hat{\boldsymbol{y}}_{p}^{u})),l_{p})\}_{p=1}^{N}$, where $\text{Vect}^{-1}(\cdot)$ is the inverse operation of $\text{Vect}(\cdot)$. To accelerate MGP Imputation, we adopt the black box matrix-matrix multiplication framework [46] to integrate the modified preconditioned conjugate gradient (mPCG) with GPU acceleration into the data imputation workflow. Please refer to Appendix VI-C for the detailed procedure for GPU acceleration of MGP imputation.
III-B MUSE-Net
@@
We further propose a MUSE-Net to process the imputed longitudinal EHRs. This model is designed to not only effectively capture crucial temporal correlations in multi-variate irregular longitudinal data but also recognize missingness patterns and account for the imbalanced data issue for enhanced classification performance. As shown in Fig. 1(b) and (c), this model consists of 3 key modules: time-aware self-attention encoder, missing value masks, and multi-branching outputs.
III-B1 Time-aware Self-attention Encoder
@@
We propose to adapt the traditional self-attention encoder [35] to incorporate the information of elapsed time between consecutive records in irregular longitudinal EHRs, including three building blocks:
@@
Elapsed time-based positional encoding: The time order of the longitudinal sequence plays a crucial role in time series analysis. However, this information is often ignored in traditional attention-based encoders because it does not incorporate any recurrent or convolution operations. To address this issue, an elapsed time-based positional encoding is added into the input sequence at the beginning of the self-attention encoder as shown in Fig. 1(b):
$PE_{t_{p,i},2m}=\text{sin}(t_{p,i}/10000^{2m/M})$ (5)
$PE_{t_{p,i},2m+1}=\text{cos}(t_{p,i}/10000^{2m/M})$
@@
where $t_{p,i}$ is the time position of each observation to account for the irregular time interval, and $2m$ and $12m+$ 1 are the ($2m$)-th and ($12m+$ 1)-th variable dimensions, respectively. The sinusoid prevents positional encodings from becoming too large, introducing extra difficulties in network optimization. The embedded elapsed time-based positional features will then be combined with the imputed signals to generate the input of the self-attention module: $X_{p}=PE_{p}+\hat{Y}_{p}$, where $PE_{p}\in\mathbb{R}^{T_{p}\times M}$ is a matrix with elements defined in Eq. (III-B1).
@@
Intepretable multi-head attention: The self-attention is a mechanism to allow the network to learn dependencies in the longitudinal data. It maps a query set and a set of key-value pairs to an output. In a single-head self-attention, the key, query, and value matrices, denoted as $K\in\mathbb{R}^{T_{p}\times M}$, $Q\in\mathbb{R}^{T_{p}\times M}$, and $V\in\mathbb{R}^{T_{p}\times M}$, are computed by taking input $X_{p}$:
@@
where $\text{softmax}(\boldsymbol{x})_{i}=\frac{e^{x_{i}}}{\sum_{j}e^{x_{j}}}$ outputs the weight assigned to each element in $V$. The dot product in the softmax is scaled down by $\sqrt{M}$. This is essential to prevent the dot product values from growing too large, especially when the dimensionality of $\hat{Y}_{p}$ or $X_{p}$ is large, introducing instabilities in the training process. The self-attention mechanism enables the network to access all the information of the input with the flexibility of focusing on certain important elements over the entire sequence.
The multi-head attention module is an extension of single-head self-attention to capture different relations among multivariate time series. Specifically, multiple matrices for the keys, queries, and values are defined by applying multiple network operations, $\mathcal{F}_{K}^{l}(\cdot)$’s, $\mathcal{F}_{Q}^{l}(\cdot)$’s, and $\mathcal{F}_{V}^{l}(\cdot)$’s, to the input:
$K_{l}=\mathcal{F}_{K}^{l}(X_{p})\in\mathbb{R}^{T_{p}\times\frac{M}{h}},~{}~{}Q_{l}=\mathcal{F}_{Q}^{l}(X_{p})\in\mathbb{R}^{T_{p}\times\frac{M}{h}}$ (8)
$V_{l}=\mathcal{F}_{V}^{l}(X_{p})\in\mathbb{R}^{T_{p}\times\frac{M}{h}}$
where $h$ is the number of attention heads and $l\in\{1,\dots,h\}$. The output of each attention head is calculated by:
$\text{Output}_{l}=\text{softmax}(\frac{Q_{l}K_{l}^{T}}{\sqrt{M/h}})\times V_{l}\in\mathbb{R}^{T_{p}\times\frac{M}{h}}$ (9)
@@
This multi-head attention module enables the network to jointly attend to information from different subspaces of multivariate time series at different time points.
However, traditional multi-head attention suffers from three limitations: (1) Lack of interpretability: the concatenation of head outputs makes it difficult to analyze how each head contributes to the final decision; (2) Independent value matrices: each head has its own $V_{l}$, leading to inconsistent feature extraction across heads; (3) Extra computation overhead: maintaining separate $V_{l}$’s increases parameter complexity. To address these issues, we adopt the Interpretable Multi-Head Attention. Instead of maintaining separate value matrices $V_{l}$, Interpretable Multi-head Attention introduces a shared value matrix $V$ for all heads $V=X_{p}W^{V}$, where $W^{V}$ is a single weight matrix applied to the entire input. Additionally, the weighted key-query matrices across all heads are summed as:
$S=\sum_{l=1}^{h}\text{softmax}(\frac{Q_{l}K_{l}^{T}}{\sqrt{M/h}})$ (11)
Then, the final output is computed by multiplying the aggregated attention score with the shared value matrix $V$:
$\text{Output}_{M}=SVW^{O}$ (12)
This modification improves interpretability by allowing a global attention map $S$ to capture how different query-key interactions influence the shared value matrix. Instead of treating each head’s output separately, interpretable multi-head attention integrates them in a structured and meaningful way.
@@
III-B2 Missing Value Masks
To further account for potential discrepancies between imputed values and actual observations, the missing value masks (Fig. 1(a)) are incorporated as an additional input to the model. We define the missing value masks for patient $p$ as $\tilde{X}_{p}$ with $[\tilde{X}_{p}]_{t,m}=1$ if $[Y_{p}]_{t,m}$ is missing; otherwise $[\tilde{X}_{p}]_{t,m}=0$. The masks will be processed by an independent time-aware self-attention encoder in parallel with the imputed sequences as shown in Fig. 1(b). This missingness-aware design enables the network to capture the relationship between non-missing and missing values. As a result, the network is capable of mitigating the effect of potential errors during the MGP imputation process, thereby enhancing the overall predictive performance of the model.
III-B3 Multi-branching Outputs
The self-attention encoder serves as a powerful feature extractor from irregular longitudinal EHRs for downstream classification tasks. The classifier network, typically with a fully connected layer, is added after the feature extractor to interpret the abstracted features and make the final prediction. The imbalanced data issue incurs a more pronounced and direct negative impact on the classifier than on the feature extractor network. This is due to the fact that the classifier network directly interacts with the labels and is heavily influenced by the majority class during the training process [50]. This leads to a bias towards the majority class, visibly affecting the model’s performance. As such, careful design of the classifier network is needed to cope with the imbalanced data issue.